PERTURBATIONS IN THE GROUND ELECTRONIC STATE ROTATIONAL SPECTRUM OF TRANSITION-METAL CONTAINING MOLECULES

\underline{D. T. HALFEN}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721; R. W. FIELD, Department of Chemistry, MIT, Cambridge, MA 02139; and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

Transition-metal containing molecules, in general, have a high density of electronic states even in the energy region of the electronic ground state. Often the pure rotational spectrum of these species exhibits perturbations caused by low-lying excited states. Examples of these perturbations include irregular spin-orbit, fine structure, and hyperfine patterns. A good example of a molecule exhibiting some of these perturbations is VCl^{+}. This species has a ^{4}Σ^{−} ground state, but has unusually large spin-spin splittings. These perturbations indicate that the ground state is interacting with the excited ^{4}Π_{i} and/or ^{2}Π_{i} states. A deperturbation analysis of the rotational spectrum of VCl^{+} is currently underway, and the results of this analysis will be presented.